OpenStudy (anonymous):

Will someone please help me with the process of how to do this?... If 4000 dollars is invested in a bank account at an interest rate of 5 per cent per year, find the amount in the bank after 12 years if interest is compounded annually, quarterly, monthly, and continuously

OpenStudy (anonymous):

For compound interest:\[A = P \times (1 + \frac{ r }{ t })^{yt}\]where r is the APR, t is the number of compounding periods per year, and y is the number of years. For compounding continuously,\[A = P \times e ^{ry}\]

OpenStudy (anonymous):

So, the easiest one is for compounding annually:\[A = 4000 \times (1 + 0.05)^{12}\]

OpenStudy (anonymous):

Awesome! I just worked it out and got it right on my homework. Thank you!

OpenStudy (anonymous):

Quarterly:\[A = 4000 \times (1 + \frac{ 0.05 }{ 4 })^{(4)(12)}\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

I'll let you use quarterly as a model to figure the monthly problem. As for continuously:\[A = 4000 \times e ^{(0.05)(12)}\]

OpenStudy (anonymous):

uw!